OpenStudy (anonymous):

What is the product of the polynomials . (8x^2+7x+9)(7x^3+9x)

OpenStudy (anonymous):

distribute the equations? have you done this before

OpenStudy (anonymous):

Is it 56x^5+49x^4+135x^3+81x?

OpenStudy (anonymous):

http://prntscr.com/1en788

OpenStudy (anonymous):

It only shows that.

OpenStudy (goldphenoix):

Yoda, is your final answer, 56x^5+49x^4+135x^3+81x ?

OpenStudy (anonymous):

Yes.

OpenStudy (mathstudent55):

@yoda5657 Your first response is not correct.

OpenStudy (anonymous):

Hmm.

OpenStudy (anonymous):

no thats not correct, try adding the like terms

OpenStudy (anonymous):

56x^5+49x^4+135x^3+63x^2+81x?

OpenStudy (goldphenoix):

I thought 56x^5+49x^4+135x^3+81x was the correct answer choices.

OpenStudy (mathstudent55):

@yoda5657 The first step must always have a number of terms that is equal to the number of the terms of the two polynomials multiplied together. You are multiplying a 3-term polynomial by a 2-term polynomial. The first step of all the multiplications has to have 3 x 2 = 6 terms.

OpenStudy (mathstudent55):

The x^2 term is missing from 56x^5+49x^4+135x^3+81x

OpenStudy (anonymous):

63^2?

OpenStudy (anonymous):

63x^2?

OpenStudy (mathstudent55):

yes

OpenStudy (anonymous):

56x^5+49x^4+135x^3+63x^2+81x

OpenStudy (mathstudent55):

Correct

OpenStudy (goldphenoix):

(8x^2+7x+9)(7x^3+9x) 8x^2 * 7x^3 = 56x^5 8x^2 * 9x = 72x^3 7x * 7x^3 = 49x^4 7x * 9x = 63x^2 9 * 7x^3 = 63x^3 9 * 9x = 81x Now you would have: 56x^5 + 72x^3 + 49x^4 + 63x^2 + 63x^3 + 81x You can comine 72x^3 and 63x^3 because the exponent are the same. 72x^3 + 63x^3 = 135x^3 Now you would have: 56x^5 + 39x^4 + 135x^3 + 63x^2 + 81x. Ouch, my bad.

OpenStudy (marissalovescats):

You have to distribute every term on the left to every term on the right and then combine like terms. 8x^2 has to go to 7x^3 and 9x 7x has to go to 7x^3 and 9x 9 has to go to 7x^3 and 9x. This making the answer 56x^5+49x^4+135x^3+63x^2+81x. Good job @yoda5657